# Quasiparticle states of hexagonal BN: A van der Waals density functional study

### v1

@@ IV Results and Discussion
Unlike graphite, h-BN experimentally demonstrates five distinct stacking arrangements, with only two being thermodynamically stable and the remaining three metastable [quian2024, ooi2006]. A key factor driving interest in h-BN is its wide band gap, which is systematically underestimated in KS-DFT [ooi2006]. In contrast, GW approximations align more closely with experimental observations [blase1995]. Here we seek to compute and contrast QP band structure results of the stable h-BN forms AA’-stacked, AB-stacked, and monolayer h-BN (BN1), Fig. 1.
@@ IV.1.1 BN1
The second-row panel of AB results shows the comparison of PBE and KI-PBE. The occupied bands of KI-PBE are down-shifted relative to PBE, while the unoccupied bands are shifted higher in energy in KI-PBE. In KI-PBE, the fundamental band gap is indirect and corresponds to the transition $\text{T}_{v}\rightarrow\text{M}_{c}$ at 6.83 eV. Here $\text{T}_{v}$ is a non-symmetric $k$ point that lies on the path from $\text{G}_{v}$ to $\text{K}_{v}$, being close to $\text{K}_{v}$. This AB-system VBM corresponds (as a shifted $\text{K}_{v}$ point) to that which was previously predicted for the AA system by use of HSE+D3 [wickramaratne2018, HSE06, grimmeD3]. We also observe that degeneracy of the unoccupied bands breaks between $\text{A}_{c}$ - $\text{H}_{c}$ and $\text{L}_{c}$ - $\text{A}_{c}$ in the KI-PBE description.
@@
The second panel of Fig. 4 shows the canonical orbital density predicted by KI-CX at the CBM. Contrasting with the corresponding BN1 panel, we find that there is now a different $p_{z}$-orbital character reflecting orbital overlaps midway between the layers. For the small $\sigma$-orbital signature of the VBMs, centered on the nitrogen atoms, there are also differences. These VBM changes arise because the vdW forces provide interlayer attraction to shorten the interlayer separation from essentially decoupled to the experimental value that characterizes our AB study [rydberg03p606, rydberg2003, Berland_2015:van_waals, shukla2022]. As such, the QP CBM changes represent vdW signatures but they should be seen as indirect [rohrer11p165423], set exclusively by the structural change. In fact, use of the KI-CX and KI-PBE (not shown in Fig. 4) at the experimental structure yields almost identical details in QP CBM orbital densities.
@@ IV.1.3 AA’
In the first-row panel of $A$ results in Fig. 3, we find a similar behavior as in the AB system. The PBE and CX results show no significant differences beyond a small displacement of valence bands, but some changes emerge for the virtual bands. The fundamental gaps emerging in PBE and CX studies reflect again an indirect $\text{T}_{v}\rightarrow\text{M}_{c}$ transition [wickramaratne2018], at 4.28 eV and 4.24 eV, respectively. We find that the VBM is at the same non-symmetric $k$ point $\text{T}_{v}$ for the AA system as for the AB system.
@@
Meanwhile, for the conduction-type QP states (unoccupied orbitals), we find that along the full path from $\text{G}_{c}$ to $\text{A}_{c}$, the use of KI-CX produces a general shift (as does KI-PBE), resembling what is documented above for the AB system. Nevertheless, the switch from CX to KI-CX also causes an enhanced lowering of the conduction-type QP energies close to $\text{M}_{c}$. The KI-CX predicts (like KI-PBE) the transition $\text{T}_{v}\rightarrow\text{M}_{c}$ [wickramaratne2018] as the fundamental gap, now at 6.58 eV.
@@
Nevertheless that are some differences. In the upper Brillouin-zone region (tracing the set of ‘A-H-L-A’ high-symmetry $k$ points of the BZ), we observe that KI-CX predicts a small but systematic and visible lowering of QP energies compared with KI-PBE. The overall lowering is expected since the reference value of the potential (by which we measure QP energy positions) is affected by the actual electron-density variation $\rho(\mathbf{r})$. While the CX and PBE densities are indeed so close that CX and PBE provide nearly identical workfunction predictions for transition metals [hyldgaard2020], there are small differences in the $\rho(\mathbf{r})$ profiles that we compute in CX (and hence in KI-CX) and in PBE (and hence in KI-PBE).
@@
Again we interpret these vdW-binding signatures as indirect QP effects because we find essentially no difference in behavior in the KI-PBE characterization of the AA system (not shown). We repeat that the PBE lacks an account of vdW and other truly nonlocal correlation effects [hyldgaard2020] and that KI-PBE cannot provide a full characterization of the QP signatures in bulk h-BN. The alignment of KI-CX and KI-PBE characterizations of QPs discussed so far only results because we consider bulk systems with the layer separations fixed at experimentally observed values. We shall return to more detailed arguments for seeing changes in the QP CBM spatial variation relative to the BN1 case, Fig. 4, as indirect vdW signatures.
@@ IV.2 Comparison with MBPT
Accuracy for BN1 QPs. Table 2 presents an overview of the quantitative results we find in the KI-CX characterization for BN1. We observe that both PBE and CX give gaps that are consistently smaller than those found in $GW_{0}$ [hunt2020]. While minimal differences exist between these functionals, PBE does yield a better agreement for the BN1 fundamental gap, at the $\text{K}_{v}\rightarrow\text{G}_{c}$ transition, while CX is in closer qualitative agreement with $GW_{0}$ in that it lifts the PBE near-degeneracy with the gap for the $\text{K}_{v}\rightarrow\text{K}_{c}$ transition. The descriptions of the QP energies can be seen as clearly improved in KI-CX.
Appendix B summarizes some challenges that we face in numerically converging these KI-CX results with respect to modeling parameters for BN1, for example, forcing us to use a higher $E_{\rm cut}^{\rm wf}$ for the conduction-band QP energies than what is required to converge the KS-DFT/KI-DFT total-energy description (or valence-band QP energies). We expect this numerical challenge arises, in part, because the screening assumption [gygi1986] that we presently need to use is not naturally set up to handle strictly two-dimensional problems. That is, future KI-CX (or KI-PBE) studies, using a better screening handling for two-dimensional systems [Sohier2017], may permit computationally cheaper QP descriptions.
Accuracy for AA QPs. Table 3 presents an overview of the accuracy that exists in PBE, CX, KI-PBE, and KI-CX predictions for the QP energies of bulk h-BN, our main focus. The assessments are made in terms of band gaps, that is, energy differences between QP transitions at the set of $k$ points that correspond to local extrema in the QP band structure. Specifically, we make quantitative comparisons of QP predictions for these band gaps by comparing with literature $GW$ results that exist for the bulk AA structure [hunt2020]. We note in passing that the $GW$ value for the AA-system fundamental gap is in close agreement with an experimental observation [arnaud2006], see Table 3. We also note that our screening approximation [gygi1986] is motivated for bulk systems and that we found these calculations robust already when picking the same $E_{\rm cut}^{\rm w}=200$ Ry value that suffices for the total-energy description, as discussed in Appendix B.
@@ IV.3 Fingerprints of vdW: QP bandstructures at CX- and PBE-predicted atomic structures
We first define general vdW fingerprints on the QP bandstructure and recall the logic of sorting these into what we call direct and indirect features, for example, as discussed in Ref. rohrer11p165423. We note that proximity of interacting fragments (in our case h-BN layers) impacts the nature [signatures, hyldgaard2020] as well as the scaling and strength of vdW forces [kleis08p205422, NanovdWScale]. We expect stronger vdW signatures to appear also in the QPs bandstructure [rohrer11p165423] as the interlayer separation shortens, something that the inclusions of the vdW-DF $E_{c}^{\rm nl}$ term itself drives. Use of CX in KI-DFT as well as in a KS-DFT structure optimization means that we 1) include vdW attraction between the layers and thus correct for the large (interlayer) underbinding that exists in a GGA like PBE [rydberg03p606, rydberg2003, shukla2022] and, in turn, that we 2) provide a consistent-KI-CX QP characterization, covering vdW impacts to the extent that they are included in CX. The correspondingly defined ‘KI-PBE(PBE)’ approach should only be used as a reference for discussions when vdW forces are important.
@@
By contrasting results for AA by KI-CX(CX) and KI-PBE(PBE), below, with those of the fixed-structure KI-CX and KI-PBE method, above, we arrive at a precise definition of direct and indirect vdW fingerprints in QP predictions. KI-CX(CX) versus KI-PBE(PBE) differences are nominally a full description of vdW and other truly nonlocal-correlation fingerprints on the QP nature. The same holds in practice for differences between KI-CX(CX) results for AA and KI-CX characterizations of BN1 (because a switch from PBE to CX has little effect on the description of intralayer bonding). Meanwhile, KI-CX versus KI-PBE differences are denoted as ‘direct’ vdW band-structure impacts because they reflect the same atom configuration and track the impact of introducing the truly nonlocal correlation term $E_{c}$ for what is, in practice, nearly identical electron-density variations $\rho(\mathbf{r})$; The similarity of CX and PBE predictions for $\rho(\mathbf{r})$ at a fixed structure has been validated by comparing workfunction shifts and vibrations across many types of systems [hyldgaard2020, grwahy20, frostenson2022]. We identify ‘indirect vdW signatures’ on our QP predictions, when KI-CX(CX) versus KI-PBE(PBE) differences cannot be identified as a direct vdW fingerprint.
@@
Before turning to specific interpretations, we express a general expectation that indirect vdW effects on QPs will be dominant for layered systems [rohrer11p165423]. This follows because PBE and hence KI-PBE(PBE) cannot drive a layered bulk system to any meaningful interlayer proximity as discussed both previously [rydberg03p606, rydberg2003] and as computed here: It is alone our CX prediction of structure that is close to the experimental characterizations of h-BN, see Section III. In any case, having the capability, via KI-CX(CX) studies, to predict both indirect and full vdW effects, is essential when we seek to predict QPs either independently from an existing experimental structure characterization [RanPRB16], or in the absence of sufficient structure data [rohrer11p165423, frostenson2022, frostenson2024, frostenson2024b], or even ahead of actual material synthesis [Sofo2007, rohrer11p165423].
@@
Contrasting the KI-CX(CX) and KI-PBE(PBE) QP results, we first focus on the valence bands along the symmetry path $\text{G}_{v}\to\text{K}_{v}\to\text{M}_{v}\to\text{G}_{v}$. At a general level, we find that use of KI-CX(CX) as our QP predictor lifts the near degeneracy found in KI-PBE(PBE) for the two top-most valence bands around the ‘K’, placing the VBM at the non-symmetric $k$-point $\text{T}_{v}$. We observe that this KI-CX(CX) VBM shift, $\text{K}_{v}\to\text{T}_{v}$, is in qualitative agreement with both KI-CX and the KI-PBE predictions obtained at the experimental value for the $c$ lattice constant, Fig. 3, lower-right panel. That is, the change is primarily a consequence of structure and should be considered as an indirect effect of including vdW interactions. We add that this VBM shift is also found when dispersion-corrected HSE-D3 [HSE03, HSE06, grimmeD3] is used to predict the band structure of the $A$ structure [wickramaratne2018].
The shift to $\text{T}_{v}$ of the QP VBM for the $A$ structure (bottom section of Table 3) is caused by a general type of nonlocal-correlation (or beyond-pure-vdW) effect that we call incipient covalency [langreth2009, Berland_2015:van_waals]. This VBM shift also arises for the AB structure, middle section of Table 3 and exists in bulk h-BN because the VBM orbital have a pronounced $p_{z}$ character [wickramaratne2018], see Fig. 4. The fact that these occupied orbitals extend into the inter-layer regions means on the one hand that they enhance the vdW attraction [hyldgaard2020] and on the other hand that such nonlocal-correlation effects drive a reduction of the interlayer separation (to close to the experimental $c$ lattice constant of bulk h-BN) compared with descriptions by semilocal functionals [rydberg03p606, rydberg2003], e.g. PBE. The overall implication is that the DFT inclusion of the nonlocal-correlation term $E_{c}^{\rm nl}$ (sometimes equated as simply vdW forces) also – and in fact mainly – causes indirect QP signatures [rohrer11p165423], for example, a small hybridization of slightly overlapping $p_{z}$ orbitals originating from different layers.
@@
The major differences between KI-CX(CX) and KI-PBE(PBE) predictions arise for the QP conduction bands, Fig. 6. For example, KI-CX(CX) predicts a $\sim 3$ eV larger dispersion in the BZ region near G (for the lowest-lying conduction bands). We find that KI-CX(CX) predicts a CBM at the $\text{M}_{c}$ symmetry point, while the KI-PBE(PBE) predictor places it near G. The difference arises because KI-CX(CX) can be seen as raising the conduction-band QP values at the $\text{G}_{c}$ point, making KI-CX(CX) consistent with $GW$ on the CBM [hunt2020]. We can interpret these effects as again being mostly indirect vdW signatures. Nevertheless, even when we compared the KI-CX and KI-PBE at the same (experimental) structure, we did find some differences for the energy-variation in QP conduction-band predictions.

### v2

@@ IV Results and Discussion
Unlike graphite, h-BN experimentally demonstrates five distinct stacking arrangements, with only two being thermodynamically stable and the remaining three metastable [128, 129]. A key factor driving interest in h-BN is its wide band gap, which is systematically underestimated in KS-DFT [129]. In contrast, GW approximations align more closely with experimental observations [130]. Here we seek to compute and contrast QP band structure results of the stable h-BN forms AA’-stacked, AB-stacked, and monolayer h-BN (BN1), Fig. 1.
@@ IV.1.1 BN1
The second-row panel of AB results shows the comparison of PBE and KI-PBE. The occupied bands of KI-PBE are down-shifted relative to PBE, while the unoccupied bands are shifted higher in energy in KI-PBE. In KI-PBE, the fundamental band gap is indirect and corresponds to the transition $\text{T}_{v}\rightarrow\text{M}_{c}$ at 6.83 eV. Here $\text{T}_{v}$ is a non-symmetric $k$ point that lies on the path from $\text{G}_{v}$ to $\text{K}_{v}$, being close to $\text{K}_{v}$. This AB-system VBM corresponds (as a shifted $\text{K}_{v}$ point) to that which was previously predicted for the AA system by use of HSE+D3 [75, 101, 132]. We also observe that degeneracy of the unoccupied bands breaks between $\text{A}_{c}$ - $\text{H}_{c}$ and $\text{L}_{c}$ - $\text{A}_{c}$ in the KI-PBE description.
@@
The second panel of Fig. 4 shows the canonical orbital density predicted by KI-CX at the CBM. Contrasting with the corresponding BN1 panel, we find that there is now a different $p_{z}$-orbital character reflecting orbital overlaps midway between the layers. For the small $\sigma$-orbital signature of the VBMs, centered on the nitrogen atoms, there are also differences. These VBM changes arise because the vdW forces provide interlayer attraction to shorten the interlayer separation from essentially decoupled to the experimental value that characterizes our AB study [123, 69, 72, 74]. As such, the QP CBM changes represent vdW signatures but they should be seen as indirect [124], set exclusively by the structural change. In fact, use of the KI-CX and KI-PBE (not shown in Fig. 4) at the experimental structure yields almost identical details in QP CBM orbital densities.
@@ IV.1.3 AA’
In the first-row panel of $A$ results in Fig. 3, we find a similar behavior as in the AB system. The PBE and CX results show no significant differences beyond a small displacement of valence bands, but some changes emerge for the virtual bands. The fundamental gaps emerging in PBE and CX studies reflect again an indirect $\text{T}_{v}\rightarrow\text{M}_{c}$ transition [75], at 4.28 eV and 4.24 eV, respectively. We find that the VBM is at the same non-symmetric $k$ point $\text{T}_{v}$ for the AA system as for the AB system.
@@
Meanwhile, for the conduction-type QP states (unoccupied orbitals), we find that along the full path from $\text{G}_{c}$ to $\text{A}_{c}$, the use of KI-CX produces a general shift (as does KI-PBE), resembling what is documented above for the AB system. Nevertheless, the switch from CX to KI-CX also causes an enhanced lowering of the conduction-type QP energies close to $\text{M}_{c}$. The KI-CX predicts (like KI-PBE) the transition $\text{T}_{v}\rightarrow\text{M}_{c}$ [75] as the fundamental gap, now at 6.58 eV.
@@
Nevertheless that are some differences. In the upper Brillouin-zone region (tracing the set of ‘A-H-L-A’ high-symmetry $k$ points of the BZ), we observe that KI-CX predicts a small but systematic and visible lowering of QP energies compared with KI-PBE. The overall lowering is expected since the reference value of the potential (by which we measure QP energy positions) is affected by the actual electron-density variation $\rho(\mathbf{r})$. While the CX and PBE densities are indeed so close that CX and PBE provide nearly identical workfunction predictions for transition metals [35], there are small differences in the $\rho(\mathbf{r})$ profiles that we compute in CX (and hence in KI-CX) and in PBE (and hence in KI-PBE).
@@
Again we interpret these vdW-binding signatures as indirect QP effects because we find essentially no difference in behavior in the KI-PBE characterization of the AA system (not shown). We repeat that the PBE lacks an account of vdW and other truly nonlocal correlation effects [35] and that KI-PBE cannot provide a full characterization of the QP signatures in bulk h-BN. The alignment of KI-CX and KI-PBE characterizations of QPs discussed so far only results because we consider bulk systems with the layer separations fixed at experimentally observed values. We shall return to more detailed arguments for seeing changes in the QP CBM spatial variation relative to the BN1 case, Fig. 4, as indirect vdW signatures.
@@ IV.2 Comparison with MBPT
Accuracy for BN1 QPs. Table 2 presents an overview of the quantitative results we find in the KI-CX characterization for BN1. We observe that both PBE and CX give gaps that are consistently smaller than those found in $GW_{0}$ [76]. While minimal differences exist between these functionals, PBE does yield a better agreement for the BN1 fundamental gap, at the $\text{K}_{v}\rightarrow\text{G}_{c}$ transition, while CX is in closer qualitative agreement with $GW_{0}$ in that it lifts the PBE near-degeneracy with the gap for the $\text{K}_{v}\rightarrow\text{K}_{c}$ transition. The descriptions of the QP energies can be seen as clearly improved in KI-CX.
Appendix B summarizes some challenges that we face in numerically converging these KI-CX results with respect to modeling parameters for BN1, for example, forcing us to use a higher $E_{\rm cut}^{\rm wf}$ for the conduction-band QP energies than what is required to converge the KS-DFT/KI-DFT total-energy description (or valence-band QP energies). We expect this numerical challenge arises, in part, because the screening assumption [125] that we presently need to use is not naturally set up to handle strictly two-dimensional problems. That is, future KI-CX (or KI-PBE) studies, using a better screening handling for two-dimensional systems [133], may permit computationally cheaper QP descriptions.
Accuracy for AA QPs. Table 3 presents an overview of the accuracy that exists in PBE, CX, KI-PBE, and KI-CX predictions for the QP energies of bulk h-BN, our main focus. The assessments are made in terms of band gaps, that is, energy differences between QP transitions at the set of $k$ points that correspond to local extrema in the QP band structure. Specifically, we make quantitative comparisons of QP predictions for these band gaps by comparing with literature $GW$ results that exist for the bulk AA structure [76]. We note in passing that the $GW$ value for the AA-system fundamental gap is in close agreement with an experimental observation [131], see Table 3. We also note that our screening approximation [125] is motivated for bulk systems and that we found these calculations robust already when picking the same $E_{\rm cut}^{\rm w}=200$ Ry value that suffices for the total-energy description, as discussed in Appendix B.
@@ IV.3 Fingerprints of vdW: QP bandstructures at CX- and PBE-predicted atomic structures
We first define general vdW fingerprints on the QP bandstructure and recall the logic of sorting these into what we call direct and indirect features, for example, as discussed in Ref. 124. We note that proximity of interacting fragments (in our case h-BN layers) impacts the nature [113, 35] as well as the scaling and strength of vdW forces [134, 135]. We expect stronger vdW signatures to appear also in the QPs bandstructure [124] as the interlayer separation shortens, something that the inclusions of the vdW-DF $E_{c}^{\rm nl}$ term itself drives. Use of CX in KI-DFT as well as in a KS-DFT structure optimization means that we 1) include vdW attraction between the layers and thus correct for the large (interlayer) underbinding that exists in a GGA like PBE [123, 69, 74] and, in turn, that we 2) provide a consistent-KI-CX QP characterization, covering vdW impacts to the extent that they are included in CX. The correspondingly defined ‘KI-PBE(PBE)’ approach should only be used as a reference for discussions when vdW forces are important.
@@
By contrasting results for AA by KI-CX(CX) and KI-PBE(PBE), below, with those of the fixed-structure KI-CX and KI-PBE method, above, we arrive at a precise definition of direct and indirect vdW fingerprints in QP predictions. KI-CX(CX) versus KI-PBE(PBE) differences are nominally a full description of vdW and other truly nonlocal-correlation fingerprints on the QP nature. The same holds in practice for differences between KI-CX(CX) results for AA and KI-CX characterizations of BN1 (because a switch from PBE to CX has little effect on the description of intralayer bonding). Meanwhile, KI-CX versus KI-PBE differences are denoted as ‘direct’ vdW band-structure impacts because they reflect the same atom configuration and track the impact of introducing the truly nonlocal correlation term $E_{c}$ for what is, in practice, nearly identical electron-density variations $\rho(\mathbf{r})$; The similarity of CX and PBE predictions for $\rho(\mathbf{r})$ at a fixed structure has been validated by comparing workfunction shifts and vibrations across many types of systems [35, 51, 102]. We identify ‘indirect vdW signatures’ on our QP predictions, when KI-CX(CX) versus KI-PBE(PBE) differences cannot be identified as a direct vdW fingerprint.
@@
Before turning to specific interpretations, we express a general expectation that indirect vdW effects on QPs will be dominant for layered systems [124]. This follows because PBE and hence KI-PBE(PBE) cannot drive a layered bulk system to any meaningful interlayer proximity as discussed both previously [123, 69] and as computed here: It is alone our CX prediction of structure that is close to the experimental characterizations of h-BN, see Section III. In any case, having the capability, via KI-CX(CX) studies, to predict both indirect and full vdW effects, is essential when we seek to predict QPs either independently from an existing experimental structure characterization [3], or in the absence of sufficient structure data [124, 102, 103, 136], or even ahead of actual material synthesis [137, 124].
@@
Contrasting the KI-CX(CX) and KI-PBE(PBE) QP results, we first focus on the valence bands along the symmetry path $\text{G}_{v}\to\text{K}_{v}\to\text{M}_{v}\to\text{G}_{v}$. At a general level, we find that use of KI-CX(CX) as our QP predictor lifts the near degeneracy found in KI-PBE(PBE) for the two top-most valence bands around the ‘K’, placing the VBM at the non-symmetric $k$-point $\text{T}_{v}$. We observe that this KI-CX(CX) VBM shift, $\text{K}_{v}\to\text{T}_{v}$, is in qualitative agreement with both KI-CX and the KI-PBE predictions obtained at the experimental value for the $c$ lattice constant, Fig. 3, lower-right panel. That is, the change is primarily a consequence of structure and should be considered as an indirect effect of including vdW interactions. We add that this VBM shift is also found when dispersion-corrected HSE-D3 [100, 101, 132] is used to predict the band structure of the $A$ structure [75].
The shift to $\text{T}_{v}$ of the QP VBM for the $A$ structure (bottom section of Table 3) is caused by a general type of nonlocal-correlation (or beyond-pure-vdW) effect that we call incipient covalency [138, 72]. This VBM shift also arises for the AB structure, middle section of Table 3 and exists in bulk h-BN because the VBM orbital have a pronounced $p_{z}$ character [75], see Fig. 4. The fact that these occupied orbitals extend into the inter-layer regions means on the one hand that they enhance the vdW attraction [35] and on the other hand that such nonlocal-correlation effects drive a reduction of the interlayer separation (to close to the experimental $c$ lattice constant of bulk h-BN) compared with descriptions by semilocal functionals [123, 69], e.g. PBE. The overall implication is that the DFT inclusion of the nonlocal-correlation term $E_{c}^{\rm nl}$ (sometimes equated as simply vdW forces) also – and in fact mainly – causes indirect QP signatures [124], for example, a small hybridization of slightly overlapping $p_{z}$ orbitals originating from different layers.
@@
The major differences between KI-CX(CX) and KI-PBE(PBE) predictions arise for the QP conduction bands, Fig. 6. For example, KI-CX(CX) predicts a $\sim 3$ eV larger dispersion in the BZ region near G (for the lowest-lying conduction bands). We find that KI-CX(CX) predicts a CBM at the $\text{M}_{c}$ symmetry point, while the KI-PBE(PBE) predictor places it near G. The difference arises because KI-CX(CX) can be seen as raising the conduction-band QP values at the $\text{G}_{c}$ point, making KI-CX(CX) consistent with $GW$ on the CBM [76]. We can interpret these effects as again being mostly indirect vdW signatures. Nevertheless, even when we compared the KI-CX and KI-PBE at the same (experimental) structure, we did find some differences for the energy-variation in QP conduction-band predictions.